OpenStudy (anonymous):

The two factors of x3 – 1331 are.. (x - 11) and (x^2 + 11x + 121)?

OpenStudy (anonymous):

Yep!

OpenStudy (anonymous):

How to check: x(x^2 + 11x + 121) - 11(x^2 + 11x + 121)

OpenStudy (anonymous):

x^3 -1331=0 x=11 so 11 is a factor of x3 – 1331 Do a long division, (x^3 -1331)/(x-11)=x^2 +11x+121 so they are the factors of x^3 -1331